# Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty

### v1

@@ 3 KKT characterization
    If $\lambda=0$, then $\mu_{ij,k}=\eta_{ji,k}>0$, which implies $x_{ij,k}=0$ and $x_{ji,k}=y_{ji,k}$ for all $i,j,k$. This is only possible if $x_{ij,k}=x_{ji,k}=y_{ji,k}=0$, a situation that will be excluded in the remainder of the analysis.
@@
    $p^{i}_{k}>p^{j}_{k}$: then
    $\eta_{ij,k}+\eta_{ji,k}>(\mu_{ij,k}+\mu_{ji,k})\,.$
    The possible ways (or combination) are
    1. (a)
      $\eta_{ij,k}>\mu_{ji,k}\geqslant 0$ and $\mu_{ij,k}=0$: $x_{ij,k}=y_{ij,k}$.
    2. (b)
      $\eta_{ji,k}>\mu_{ij,k}\geqslant 0$ and $\mu_{ji,k}=0$: $x_{ji,k}=y_{ji,k}$.
  3. 3.
    $p^{j}_{k}>p^{i}_{k}$: then
    $\eta_{ij,k}+\eta_{ji,k}<(\mu_{ij,k}+\mu_{ji,k})\,.$
    The possible ways (or combination) are
    1. (a)
      $0\leqslant\eta_{ji,k}<\mu_{ij,k}$ and $\eta_{ij,k}=0$: $x_{ij,k}=0$.
    2. (b)
      $0\leqslant\eta_{ij,k}<\mu_{ji,k}$ and $\eta_{ji,k}=0$: $x_{ji,k}=0$.
  4. 4.
    $p^{j}_{k}=p^{i}_{k}\geqslant 0$: then
    $\eta_{ij,k}+\eta_{ji,k}=(\mu_{ij,k}+\mu_{ji,k})\,.$
    $\eta_{ij,k}$, $\eta_{ji,k}$, $\mu_{ij,k}$ and $\mu_{ji,k}$ cannot all be positive (a situation previously isolated). The possible ways (or combinations) are
    1. (a)
      $\mu_{ij,k}=\eta_{ji,k}>0$ and $\eta_{ij,k}=\mu_{ji,k}=0$: $x_{ij,k}=0$ and $x_{ji,k}=y_{ji,k}$.
    2. (b)
      $\mu_{ji,k}=\eta_{ij,k}>0$ and $\eta_{ji,k}=\mu_{ij,k}=0$: $x_{ji,k}=0$ and $x_{ij,k}=y_{ij,k}$.
    3. (c)
      $\eta_{ij,k}=\eta_{ji,k}=\mu_{ij,k}=\mu_{ji,k}=0$ and the constraints on $X$ are not satiated.
Now let us analyze a Nash equilibrium situation: Consider the market of good $k$. Given a customer $i$, let designate $\mathbb{E}^{+}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}>p^{i}_{k}\}$, $\mathbb{E}^{-}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}<p^{i}_{k}\}$ and $\mathbb{E}^{=}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}=p^{i}_{k}\}$.
  1. 1.
    For $j\in\mathbb{E}^{-}_{k}$ (resp. $j\in\mathbb{E}^{+}_{k}$), the KKT condition for both customers implies one of the three situations:
    1. (a)
      $x^{i}_{ij,k}=x^{j}_{ij,k}=0$ and $x^{i}_{ji,k}=x^{j}_{ji,k}$.
    2. (b)
      $x^{i}_{ij,k}=x^{j}_{ij,k}$ and $x^{i}_{ji,k}=x^{j}_{ji,k}=0$.
    3. (c)
      $x^{i}_{ij,k}=x^{j}_{ij,k}=0=x^{j}_{ji,k}=x^{i}_{ji,k}$.
  2. 2.
    For $j\in\mathbb{E}^{=}_{k}$, the Nash equilibrium condition implies:
    1. (a)
      $0\leqslant x^{i}_{ij,k}=x^{j}_{ij,k}$ and $0\leqslant x^{i}_{ji,k}=x^{j}_{ji,k}$.
Remark 3.1.
To avoid the indeterminacy caused by the situation $\mathbb{E}^{=}_{k}$, particularly the occurrence of unnecessary transactions in both directions ($0<x^{i}_{ij,k}$ and $0<x^{i}_{ji,k}$) when prices are equal, one could impose the condition $x^{i}_{ij}\cdot x^{i}_{ji}=0$ for all $i,j$. This condition ensures that no simultaneous flows occur in opposite directions and will be implicitly assumed when considering the joint feasible set to guarantee boundedness.
4 Convergence
Next, we present three convergence processes toward Pareto optimal and Nash–Pareto states.
@@ 4.1 Gradient direction
A classical proof of the projected supergradient descend convergence could be found in [Sho98]. The convergence point $\left(p^{\star},X^{\star}\right)=\left(p^{1,\star},X^{1,\star},\ldots,p^{n,\star},X^{n,\star}\right)$ satisfies
@@ 7 Topological Influences on Consumer Choice and Market Coordination
In the following, we address the problem of demand indeterminacy by analyzing consumer preferences within an exchange network. Our focus is on exchange processes structured by network interactions: each agent $i$ may hold intrinsic preferences over goods, as captured by classical demand theory (see, for instance, [MCWG95]), while also developing affinities toward specific trading partners. These affinities—shaped by factors such as marketing, customer loyalty, market positioning, or interpersonal relationships—can lead agents to favor transactions with certain peers over others. We proceed to formalize and examine this layered preference structure and explore its implications for decentralized exchange.
@@
In practice, the function $v$ can be made more explicit. For example, one could consider $v$ to be some measure of centrality in the network, as there are many possibilities. From a trade perspective, one could imagine that the seller, depending on their position in the market, could have different preferences. For instance, a large seller might seek to gain more influence in the market and thus maximize their spectral influence (see [Ria24]), while another agent might aim to maximize their prestige by forming relationships with larger sellers and thus maximize their eigenvector centrality. For more details on centrality measures, consult [BJT23]. Marketing and social determinants can also influence the form of $v$.
@@ 8 Numerical simulation: Exchange with indivisible goods
This numerical example is crucial for understanding the price vector $p$. In our model, price is subjective: a null price $p^{i}_{k}$ for agent $i$ indicates that agent $i$ is not involved in transactions of good $k$. In other words, an agent assigns a price only to goods they intend to trade. This is natural, as price is a market-related concept determined by supply and demand forces. It should not be confused with intrinsic value. [^2] The distinction between use value and exchange value can be traced back to Adam Smith [Smi76]. Furthermore, the topology plays a crucial role in shaping equilibria, influencing welfare losses, and highlighting imperfections caused by intermediaries.
@@ 10.1 Equilibrium with money
The realistic nature of transactions makes the Clower cash-in-advance hypothesis unnecessary [Clo67], since each individual only demands goods that they can afford, and every transaction is equivalent to a monetary flow. When analyzing the monetary flows, the customer satisfies the following transaction balance condition
@@ 11.1 Transaction equilibrium in a simple financial market
where $S^{i_{s}}_{t}=\sum_{j\neq i_{s}}\left(p^{i_{s}}(t)\cdot x^{i_{s}}_{i_{s}j}(t)-p^{j}(t)\cdot x^{i_{s}}_{ji_{s}}(t)\right)+(1-\delta)S^{i_{s}}_{t-1}$ denotes the producer’s stock of commodities at time $t$; $\delta$ is the depreciation rate of the stock; $\mathscr{S}^{i_{s}}_{t}$ is a closed, bounded subset of $\mathbb{R}_{+}^{L}$ representing storage capacity constraints; and $\tau=(\tau_{1},\dots,\tau_{L})$ is a vector of time requirements for producing the outputs $y^{i_{s}}(t+\tau)$, with time being treated as a production factor. The time gap between production and exchange results in an intertemporal transfer [^3] A similar idea was introduced by Keynes under the notion of effective demand in his book [Key36].. The production-possibility set $\mathscr{Y}^{i_{s}}_{t}$ of producer $i_{s}$} satisfies, at each time $t$:
@@ 12 Uncertainty
Under uncertainty, agents are assumed to form expectations about future events to guide both their current and future decisions. In contrast to the rational expectations hypothesis ([Mut61]), which is grounded in conditional expectations, we argue that expectations should instead be formed based on the mode. The rationale is that individuals are more inclined to act upon the most likely outcome rather than an average. Relying on the mean of two distinct possibilities may result in decisions that are suboptimal for both, whereas empirical evidence suggests that individuals typically focus on the most probable event as the one most worth preparing for. While situations of indecision may arise—where multiple outcomes appear equally plausible—individuals often resolve this ambiguity by relying on personal beliefs, social norms, cognitive simplification, or external influences. In practice, these mechanisms contribute to the emergence of a unimodal distribution.
@@ 12.1 The mode as an anticipation criterion
So $g(z,x)\to 0$ as $\|z\|\to\infty$. One could then apply the generalized maximum Theorem 14.4 (as proved in [FKV14] - theorem 1.2), to prove that the mode is an an upper semicontinuous map $\mathbb{M}(x)$, since $F$ is a constant set-valued map. Moreover, under the strict quasiconcavity of $g(x,\cdot)$, one could end up with a continuous function.
@@ 14 Conclusion
  • Agent anticipation is formed by the mode, incorporating personal information, beliefs, and limited memory. The use of the mean as an expectation does not accurately reflect real-world behavior. Instead, anticipation should reflect subjective judgment, bounded rationality, and memory constraints. In particular, agents are assumed to base their forecasts on the most likely outcome—captured by the mode—which aligns more closely with Keynesian notions of "animal spirits" [Key36].
@@ 14.1 Berge’s Maximum Theorem
Berge’s Maximum Theorem[Ber63]
@@ 14.2 Generalized Maximum Theorem
Compactly generated spaces[FKV14]

### v2

@@ 4 Convergence
    A gradient-based algorithm representing an exchange dynamic driven by agents’ needs and desires (utility gradient);
@@
    A non-tâtonnement adjustment process where agents trade as long as further improvement is attainable.
@@ 4.1 Gradient direction
A classical proof of the projected supergradient descend convergence could be found in [Shor, 1998]. The convergence point $\left(p^{\star},X^{\star}\right)=\left(p^{1,\star},X^{1,\star},\ldots,p^{n,\star},X^{n,\star}\right)$ satisfies
@@ 7 Topological Influences on Consumer Choice and Market Coordination
In the following, we address the problem of demand indeterminacy by analyzing consumer preferences within an exchange network. Our focus is on exchange processes structured by network interactions: each agent $i$ may hold intrinsic preferences over goods, as captured by classical demand theory (see, for instance, [Mas-Colell et al., 1995]), while also developing affinities toward specific trading partners. These affinities—shaped by factors such as marketing, customer loyalty, market positioning, or interpersonal relationships—can lead agents to favor transactions with certain peers over others. We proceed to formalize and examine this layered preference structure and explore its implications for decentralized exchange.
@@
In practice, the function $v$ can be made more explicit. For example, one could consider $v$ to be some measure of centrality in the network, as there are many possibilities. From a trade perspective, one could imagine that the seller, depending on their position in the market, could have different preferences. For instance, a large seller might seek to gain more influence in the market and thus maximize their spectral influence (see [Riane, 2024]), while another agent might aim to maximize their prestige by forming relationships with larger sellers and thus maximize their eigenvector centrality. For more details on centrality measures, consult [Bloch et al., 2023]. Marketing and social determinants can also influence the form of $v$.
@@ 8 Numerical simulation: Exchange with indivisible goods
This numerical example is crucial for understanding the price vector $p$. In our model, price is subjective: a null price $p^{i}_{k}$ for agent $i$ indicates that agent $i$ is not involved in transactions of good $k$. In other words, an agent assigns a price only to goods they intend to trade. This is natural, as price is a market-related concept determined by supply and demand forces. It should not be confused with intrinsic value. [^2] The distinction between use value and exchange value can be traced back to Adam Smith [Smith, 1776]. Furthermore, the topology plays a crucial role in shaping equilibria, influencing welfare losses, and highlighting imperfections caused by intermediaries.
@@ 10.1 Equilibrium with money
The realistic nature of transactions makes the Clower cash-in-advance hypothesis unnecessary [Clower, 1967], since each individual only demands goods that they can afford, and every transaction is equivalent to a monetary flow. When analyzing the monetary flows, the customer satisfies the following transaction balance condition
@@ 11.1 Transaction equilibrium in a simple financial market
where $S^{i_{s}}_{t}=\sum_{j\neq i_{s}}\left(p^{i_{s}}(t)\cdot x^{i_{s}}_{i_{s}j}(t)-p^{j}(t)\cdot x^{i_{s}}_{ji_{s}}(t)\right)+(1-\delta)S^{i_{s}}_{t-1}$ denotes the producer’s stock of commodities at time $t$; $\delta$ is the depreciation rate of the stock; $\mathscr{S}^{i_{s}}_{t}$ is a closed, bounded subset of $\mathbb{R}_{+}^{L}$ representing storage capacity constraints; and $\tau=(\tau_{1},\dots,\tau_{L})$ is a vector of time requirements for producing the outputs $y^{i_{s}}(t+\tau)$, with time being treated as a production factor. The time gap between production and exchange results in an intertemporal transfer [^3] A similar idea was introduced by Keynes under the notion of effective demand in his book [Keynes, 1936].. The production-possibility set $\mathscr{Y}^{i_{s}}_{t}$ of producer $i_{s}$} satisfies, at each time $t$:
@@ 12 Uncertainty
Under uncertainty, agents are assumed to form expectations about future events to guide both their current and future decisions. In contrast to the rational expectations hypothesis ([Muth, 1961]), which is grounded in conditional expectations, we argue that expectations should instead be formed based on the mode. The rationale is that individuals are more inclined to act upon the most likely outcome rather than an average. Relying on the mean of two distinct possibilities may result in decisions that are suboptimal for both, whereas empirical evidence suggests that individuals typically focus on the most probable event as the one most worth preparing for. While situations of indecision may arise—where multiple outcomes appear equally plausible—individuals often resolve this ambiguity by relying on personal beliefs, social norms, cognitive simplification, or external influences. In practice, these mechanisms contribute to the emergence of a unimodal distribution.
@@ 12.1 The mode as an anticipation criterion
So $g(z,x)\to 0$ as $\|z\|\to\infty$. One could then apply the generalized maximum Theorem 14.4 (as proved in [Feinberg et al., 2014] - theorem 1.2), to prove that the mode is an an upper semicontinuous map $\mathbb{M}(x)$, since $F$ is a constant set-valued map. Moreover, under the strict quasiconcavity of $g(x,\cdot)$, one could end up with a continuous function.
@@ 14 Conclusion
  • Agent anticipation is formed by the mode, incorporating personal information, beliefs, and limited memory. The use of the mean as an expectation does not accurately reflect real-world behavior. Instead, anticipation should reflect subjective judgment, bounded rationality, and memory constraints. In particular, agents are assumed to base their forecasts on the most likely outcome—captured by the mode—which aligns more closely with Keynesian notions of "animal spirits" [Keynes, 1936].
@@ 14.1 Berge’s Maximum Theorem
Berge’s Maximum Theorem[Berge, 1963]
@@ 14.2 Generalized Maximum Theorem
Compactly generated spaces[Feinberg et al., 2014]
@@
Declarations
Conflict of interest
The author declares that they have no conflict of interest.
Authors’ contributions
The author carried out the study conception and design. Material preparation, data generation and analysis were performed by the author. The manuscript was written by the author.
Funding
No funding was received for conducting this study.
Availability of data and materials
All computations were carried out with Wolfram Mathematica.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author used ChatGPT in order to improve the text. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.